# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

### v1

@@ 10 Ablation Study on Model Architecture
Spatiotemporal Attention
In the default setting, we follow the AuxFormer[41] of computing temporal and spatial attention sequentially. Results of parallel computation of temporal and spatial attention are presented in Tab. 8 C, showing worse performance.
Model Size
In Tab. 9, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results. Although reducing the model size can decrease the number of parameters, it can also lead to decreased performance. We observe that the default setting ($d$=256, $l$=3) represents the optimal balance between parameter count and performance.

### v2

@@ 12 Visualization of Attention Feature Map
13 Full Results on Human3.6M
Motion Walking Eating Smoking Discussion Direction
Mixer[bouazizi2022motionmixer] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
PGBIG[ma2022progressively] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
SPGSN[li2022skeleton] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
siMLPe[guo2023back] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
DSTD-GC[fu2023learning] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
AuxFormer[xu2023auxiliary] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
CIST-GCN[medina2024context] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
GCNext[wang2024gcnext] 8.8 38.9 5.9 35.0 5.6 36.1 8.8 63.1
ALIEN[wei2025alien] 9.5 18.7 33.8 40.2 6.6 14.6 30.5 38.1 6.4 13.9 28.8 35.7 9.6 23.7 53.8 67.2 6.6 16.8 40.5 51.4
Motion Greeting Phoning Posing Purchases Sitting
Mixer[bouazizi2022motionmixer] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
PGBIG[ma2022progressively] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
SPGSN[li2022skeleton] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
siMLPe[guo2023back] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
DSTD-GC[fu2023learning] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
AuxFormer[xu2023auxiliary] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
CIST-GCN[medina2024context] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
ALIEN[wei2025alien] 13.6 32.1 70.3 85.8 8.1 18.2 39.3 49.4 9.8 25.1 61.2 78.9 12.2 28.9 62.0 76.6 8.7 19.2 42.9 54.9
Mixer[bouazizi2022motionmixer] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
PGBIG[ma2022progressively] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
SPGSN[li2022skeleton] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 8.9 18.2 33.8 40.9
siMLPe[guo2023back] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
DSTD-GC[fu2023learning] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
AuxFormer[xu2023auxiliary] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
CIST-GCN[medina2024context] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
ALIEN[wei2025alien] 13.4 27.0 56.4 70.5 8.8 19.0 42.1 53.5 8.3 19.2 43.3 54.8 18.7 40.3 75.9 88.5 8.4 17.8 33.7 40.4